OpenStudy (anonymous):

Find the derivative: y= sin(cos(2x))

OpenStudy (abb0t):

Use the chain rule. The derivative of the outside times the derivative of the inside.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

-sin(2) * sin?

OpenStudy (abb0t):

Hence, you want to find f' of sin first, then multiply it by the derivative of cos (the inside), and so forth with the 2x.

OpenStudy (anonymous):

So, sin(2x) = -cos2 right?

OpenStudy (abb0t):

You keep the (inside) you keep, you are sort of treating it as a constant. For example: \[\sin(x-p+y^{zt}-\frac{ \sqrt{s} }{ M }) \] then f' is: \[\cos(x-p+y^{zt}-\frac{ \sqrt{s} }{ M }) \times \frac{ d }{ dx }[(x-p+y^{zt}-\frac{ \sqrt{s} }{ M })]\]

OpenStudy (anonymous):

Understood, thanks!